Hydrogen initially occupies 183 liters. If it changes to a volume of 58 L at 0° C and 760 mm Hg. If the initial temperature of the hydrogen was 77° C, what was its initial pressure of the gas in mm Hg? Round to two decimal places.


1
Expert's answer
2021-02-26T05:59:04-0500

pV=nRT.

n=pV/RT = 760 x 58/62400x273 = 0,00258758336 moles.

p=nRT/V= 0,00258758336 x 62400 x 350/183 = 308,813228 mmHg.